# TRATES: Trait-Specific Rubric-Assisted Cross-Prompt Essay Scoring

### v1

@@ Appendix H Rubric-based Assessment Questions
Table 8: The total number of LLM-generated questions from all the unique rubrics of each trait for the ASAP dataset.
@@
Table 8 shows the total number of generated questions by each LLM from the rubrics of each trait. Additionally, Table H presents examples of the assessment questions generated by the Starling LLM for each rubric. These rubrics are derived from the ASAP and ASAP++ datasets Mathias and Bhattacharyya (2018), which were utilized in our evaluation. The data indicates that Llama tends to generate the most numerous questions, followed by Starling and then Gemma.
For ELLIPSE dataset, one rubric for each trait is used to score all the essays in the dataset. Table H illustrates the number of generated questions from each rubric, and an example of the generated questions for each trait.

### v2

@@ 4 Activation ReLU ReLU, SELU, LeakyReLU, Tanh, ELU
Table 9: The number of LLM-generated questions from all the unique rubrics of each trait for the ASAP dataset.
@@
Appendix B Generic Writing-Quality Features
  1. 1.
    Length-based features, such as the number of words and sentences in the essay; these features are considered the most intuitive indicators for writing quality, which have been used extensively over the years Mathias and Bhattacharyya (2018); Chen and Li (2023),
  2. 2.
  3. 3.
    Text variations features covering the usage of part-of-speech tags (POS) and punctuation.
  4. 4.
  5. 5.
Appendix C ASAP Dataset Statistics
The ASAP++ comprises 12,978 essays written in English in response to 8 prompts. Each of these prompts has a different number of responses and a different score range. There are a total of 10 different traits; each is covered in a subset of the prompts, except the content trait which is covered in all prompts. It’s important to note that the traits of voice and style are only addressed in one prompt, making them unsuitable for cross-prompt experiments. Table 6 presents the statistics and the details of the dataset.
Appendix D ELLIPSE Dataset Statistics
Appendix E LLMs Selection
We selected three LLMs for our experiments:
  1. 1.
  2. 2.
  3. 3.
Appendix F Zero-shot LLM Scoring
Figure 4: LLM prompt for Zero-shot LLM direct scoring (LLM-D).
Appendix G Hyperparameters
Appendix H Rubric-based Assessment Questions
Table 9 shows the total number of generated questions by each LLM from the rubrics of each trait. Additionally, Table A presents examples of the assessment questions generated by the Starling LLM for each rubric. These rubrics are derived from the ASAP and ASAP++ datasets Mathias and Bhattacharyya (2018), which were utilized in our evaluation. The data indicates that Llama tends to generate the most numerous questions, followed by Starling and then Gemma.
For ELLIPSE dataset, one rubric for each trait is used to score all the essays in the dataset. Table A illustrates the number of generated questions from each rubric, and an example of the generated questions for each trait.